As a gasoline engine is running, an amount of gasoline containing 15,500 J of chemical potential energy is burned in 1 s. During that second, the engine does 3,100 J of work.
(a)
What is the engine's efficiency (in percent)?

(b)
The burning gasoline has a temperature of about 4,400°F (2,700 K). The waste heat from the engine flows into air at about 82°F (301 K). What is the Carnot efficiency (in percent) of a heat engine operating between these two temperatures?
1
Expert's answer
2019-04-22T11:00:34-0400

A) By definition, efficiency is η = 100% * 3100/15500 = 20%


B) In accordance with the theory of Carnot  efficiency η = 100%*(1 - Ts/Tf)  = 100%*(1 - 301/2700) = 89%.